How can i solve this problem in
solving system in equations using subistituion
can you show me how to start and finished i simpliy dont understand it
x+y=6
-x+y=4

Answer by Alan3354(69443)

x+y=6
-x+y=4
---------
Elimination is best for this system.
======
x+y=6
-x+y=4
-------- Add
2y = 10
y = 5
---
x = 1
===========
Using substitution:
----
x+y=6
-x+y=4 --> y = x+4
Sub for y in the 1st eqn
x + (x+4) = 6
2x + 4 = 6
2x = 2
x = 1
----
etc